OpenStudy (anonymous):

verify that x^4-y^4=(x-y)(x^3+x^2y+xy^2+y^3)

jhonyy9 (jhonyy9):

if you consider x*4 -y*4 = (x*2)*2 -(y*2)*2 =(x*2 -y*2)(x*2 +y*2)=(x-y)(x+y)(x*2 +y*2)=(x-y)(x*3 +xy*2 +x*2y +y*3)

jhonyy9 (jhonyy9):

hope that you can understanding

OpenStudy (anonymous):

thank you very much....

jhonyy9 (jhonyy9):

please you